Find the x - intercept and y - intercept of the line given by the equation
9x + 4y = -10

Answer by rfer(16322)

9x+4y=-10
9x+4*0=-10
x=-10/9
(-10/9,0)
9*0+4y=-10
4y=-10
y=-2.5
(0,-2.5)